OpenStudy (anonymous):

Solve for x (2/5) + (3/5x) = (x+5)/10 I keep getting stuck at 6 = x^2 -x

OpenStudy (anonymous):

its quadratic solve it by dharacharyas formula

OpenStudy (anonymous):

ans is 13 0r 12

OpenStudy (anonymous):

Well, that isn't any of the answers given. We are supposed to be using something similar to how we were simplifying rational expressions. When i follow those steps I find the common denominator, and the corresponding numerators to get: (4x/10x) + (6/10x) = ((x^2 + 5x)/10x) [I distributed the x from x(x+5) that i got from the numerator after finding the corresponding values with the denominator] At this point, my lesson said I should be able to just remove the common denominators and solve for x like normal, but it just isn't working.

OpenStudy (anonymous):

4x + 6 = x^2 +5x x^2 + x - 6 = 0 X^2 + 3x -2x +6= 0 x ( x+3) -2 ( x+3) = 0 (x-2) (x+3) =0 x =2 or -3

OpenStudy (anonymous):

Oh! That makes perfect sense, and is an answer choice! I see what I did wrong now, I forgot that I was putting the equation equal to 0 Then it was just a matter of factoring! Thank you for the help Meinme :)

OpenStudy (anonymous):

mention not....